A 110 board is cut into two pieces. One piece is three times the length of the other. find the lengths of the two pieces?
I tried x+3x=110
4x=110
x=110/4
but then I got stuck

Answer by DrBeeee(684)

x = 110/4 or
x = 27.5
Is (27.5 + 3*27.5 = 110)?
Is (27.5 + 82.5 = 110)?
Is ( 110 = 110)? Yes
Answer: one piece is 27.5 units long and the other is 82.5 units long.